40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 22 18:15:21 2010, kiwamu, Update, Locking, Green era: found a green beat note finally green_beatnote.jpg
    Reply  Fri Oct 22 21:28:27 2010, Aidan, Update, Locking, Green era: found a green beat note finally 
       Reply  Sun Oct 24 18:06:35 2010, kiwamu, Summary, Locking, setup for green beat green_beat_note.png
          Reply  Mon Oct 25 02:14:38 2010, Koji, Summary, Locking, setup for green beat 
Message ID: 3774     Entry time: Mon Oct 25 02:14:38 2010     In reply to: 3771
Author: Koji 
Type: Summary 
Category: Locking 
Subject: setup for green beat 

- What is the actual photocurrent for the beam1 and beam2? We don't care how much power do you have before the BS, but care how much current do you have on the PD.

- How much is the visibility? There is mismatching of the beams. i.e. The beam diameter looked quite different. Also the beams are not TEM00 but have fringes probably comes from the TT mirrors. You maybe able to measure the visibility by the DC output, making the beat freq go through df=0 slowly.

- What is the measured gain of the RF amp? Does it include the voltage division by the output/input impedance?

---------------------

 The signal level of the observed peak was -48dBm.

However I was expecting it is like -28dBm with some ideal assumptions.

There may be a 20dB unknown loss which sounds big to me.

I was assuming the parameters are like:

           A = 0.39 [A/W]   (assuming 90% quantum efficiency at 532nm)

           Z = 240 [V/A]  

           P1 = 17 uW  (measured by Newport power meter)

           P2 = 30 uW (measured by Newport power meter)

           G_RF = 23 dB

ELOG V3.1.3-